### v2

# Enhancing Interaction Modeling with Agent Selection and Physical Coefficient for Trajectory Prediction

Abstract
A thorough understanding of the interaction between the target agent and surrounding agents is a prerequisite for accurate trajectory prediction. Although many methods have been explored, they all assign correlation coefficients to surrounding agents in a purely learning-based manner. In this study, we present ASPILin, which manually selects interacting agents and calculates their correlations instead of attention scores. Surprisingly, these simple modifications can significantly improve prediction performance and substantially reduce computational costs. Additionally, ASPILin models the interacting agents at each past time step separately, rather than only modeling the interacting agents at the current time step. This clarifies the causal chain of the target agent’s historical trajectory and helps the model better understand dynamic interactions. We intentionally simplified our model in other aspects, such as map encoding. Remarkably, experiments conducted on the INTERACTION, highD, and CitySim datasets demonstrate that our method is efficient and straightforward, outperforming other state-of-the-art methods.
I Introduction
The ability to accurately forecast the trajectories of human-driven vehicles and pedestrians sharing the environment with autonomous vehicles is paramount within autonomous driving. Such precise trajectory predictions are indispensable for downstream intelligent planning systems to make informed decisions, thereby improving autonomous driving operations’ safety, comfort, and efficiency. However, due to the inherent uncertainty and the multi-modal nature of driving behaviors, vehicle trajectory prediction against an urban setting presents significant challenges that include, but are not limited to, spatio-temporal modeling of historical trajectories[1], interaction modeling[2, 3], environmental description[4], kinematic constraints[3, 5], and real-time inference[6].
Recent studies[7, 2, 3, 8, 9] focus on modeling interactions between agents, as it is a crucial element in autonomous driving. For instance, HiVT[7] and QCNet[9] leverage the attention mechanism to model agent-agent interactions, which can implicitly select significant nearby agents for the target agent. The larger attention weights indicate the importance of these agents. Although the results of these studies demonstrate the superiority of their interaction modeling, they still have some shortcomings:
(i) Previous methods tend to take all surrounding agents as input to the interaction module. However, human attention capacity is limited. In dynamic environments, one person can focus on at most 5 agents at a time[10], meaning a huge amount of irrelevant agents are also input into the model, which significantly increases the model’s inference latency and may also reduce prediction performance, especially in interaction-rich scenarios.
(ii) The interpretability of interaction encoding is low. Attention mechanism[1] and Graph Neural Network (GNN)[11] are popular in numerous interaction-aware methods. Fundamentally, these methods do not elucidate the underlying decision logic or cognitive processes. Although recent methods[3, 5, 12] use a combination of physics- and learning-based approaches, their interaction encoding remains purely learning-based.
Figure 1: Left: We first predict each agent’s future lane. Then, combine it with their current lane and direction relative to the target agent to select interacting agents further. Right: Classical agent selection method, which takes all agents within a defined range as model input.
To this end, we present ASPILin, a simple Conditional Variational Autoencoder (CVAE) framework that utilizes reparameterization to predict future trajectories. For the interaction module, firstly, we introduce a more refined scheme for agent selection as illustrated in Fig. 1. However, a new challenge arises as well: if we, like most approaches[13, 12, 11], only model the interacting agents of the target agent at the current moment, early interacting agents may be ignored, leading to an incomplete causal chain of the entire trajectory. Thus, we perform agent selection at each historical moment and retain only their states at that moment, then concatenate the states from all moments to create a complete interaction representation. Lastly, we derive a novel correlation coefficient by estimating agents’ future nearest position, the time needed to reach that, and their current distance. All correlation coefficients are then normalized across the spatio-temporal dimensions to replace attention scores and are integrated into the Transformer framework. We deliberately used simple network structures to simplify our model for the other two modules (historical trajectory encoding and map encoding). Comparative experiments on the INTERACTION[14] and highD[15] datasets demonstrate that ASPILin is highly competitive with other state-of-the-art methods. More importantly, ablation studies on the INTERACTION and CitySim[16] datasets indicate that our improvements to the interaction module achieved better prediction performance with lower inference latency. We also highlight that these improvement strategies can be easily incorporated into other models, particularly the agent selection strategy.
II Related Work
II-A Interaction-Aware Trajectory Prediction
Existing interaction-aware methods can be further explored from the perspectives of agent selection and interaction encoding.
For interacting agent selection, some methods[17, 18] directly model interactions with all agents within the scene and simultaneously predict the trajectories of multiple target agents. By contrast, setting a range threshold[4, 3, 13, 1, 2, 11, 19] or limiting the maximum number of neighbors[8] permits modeling of the target agent’s local context, which aligns more closely with the needs of single-agent prediction[4, 13, 1]. Moreover, many works[1, 8, 7] conduct joint predictions by initially capturing local interactions before modeling global interactions, enabling the model to extend from single-agent prediction to multi-agent prediction. Most of these methods only model the interacting agents of the target agent at the current moment. For multi-step prediction methods[3] and vectorized representation methods[4, 7], they must model the interactions for each historical timestep.
For interaction encoding, all studies use purely learning-based approaches. However, in other respects, recent studies combine physics- and learning-based approaches, offering insights into improving model performance. SSP-ASP[5] and ITRA[3] limit motion learning to an action space grounded in acceleration and steering angles, subsequently deducing future trajectories via a kinematic model. M2I[12] classifies a pair of agents as influencer and reactor by calculating the closest value of their ground-truth trajectories and the time required to reach the nearest point at the training stage, followed by the sequential generation of their future trajectories via a marginal predictor and a conditional predictor, respectively.
II-B Multi-modal Trajectory Prediction
The future trajectory of vehicles inherently exhibits multimodality, given the uncertainty of intentions. To tackle this challenge, one widely used approach involves modeling the output as a probability distribution of future trajectories via regression[7]. Usually, it introduces a cross-entropy loss function for mode classification to avoid mode collapse. Some methods use more explicit classification representations to make multi-modal trajectories closer to reality. TNT[20] samples anchor points from the roadmap and then generates trajectories based on these anchors. SSL-Lanes[19] classifies the maneuvers of each agent and trains the model in a self-supervised manner.
Other methods parameterize the distribution of future trajectories[21], such as Gaussian Mixture Models (GMM) or samples within a latent space, and generate predictions through mapping. Regarding the latter, Generative Adversarial Networks (GANs)[22], Conditional Variational Autoencoders (CVAEs)[1, 23], and diffusion model[6] are the most popular models. A common drawback of generative models is the need for extensive data to support training. Moreover, for GANs, challenges such as training difficulties and mode collapse exist. For the diffusion model, multi-step denoising leads to significant computational overhead and high inference latency. Although CVAEs face issues of insufficient diversity like GANs, their training is more stable. In this study, instead of sampling randomly from a standard normal distribution, we treat the sampler as a trainable module to prevent unrealistic trajectory outputs due to randomness.
Figure 2: The illustration of ASPILin. We focus on interaction modeling and deliberately simplify the design of other modules to prove the effectiveness of our method. A lane predictor and a new algorithm are used to select interactive agents further, and a novel physical correlation coefficient is designed to replace data-driven attention encoding. The multimodal trajectory predicion results are finally derived from a reparameterization formula.
III Methodology
III-A Problem Formulation
Single-agent trajectory prediction is designed to forecast the future trajectory of a target agent conditioned on agents’ historical states $X$ and the map information $\mathcal{M}$. To be more specific, we assume that at time $t$, there are $N$ agents (vehicles, pedestrians, cyclists) in the scene, so their historical states can be represented as ${X_{t}}=[x_{t}^{1},x_{t}^{2},\ldots,x_{t}^{N}]\in\mathbb{R}^{N\times 7}$, where $x_{t}^{n}$ is the state of agent $n$ at time $t$, including $xy$ coordinates ($p_{t}^{n}\in\mathbb{R}^{2}$), heading angle ($h_{t}^{n}\in\mathbb{R}$), speed ($v_{t}^{n}\in\mathbb{R}^{2}$), and acceleration ($a_{t}^{n}\in\mathbb{R}^{2}$). Specifically, $x_{t}^{1}$ represents the state of the target agent at time $t$. Taking into account $T_{h}$ historical observation timesteps, the overall historical state of the agents is denoted as $X=[X_{-T_{h}+1},X_{-T_{h}+2},\ldots,X_{0}]\in\mathbb{R}^{N\times T_{h}\times 7}$. Similarly, the future ground-truth trajectory of the target agent is defined as $Y=[y_{1},y_{2},\ldots,y_{T_{f}}]\in\mathbb{R}^{T_{f}\times 2}$ over $T_{f}$ timesteps, where $y_{t}$ is the $xy$ coordinates at time $t$. To forecast multi-modal future trajectories, the predicted trajectory of $K$ modes is denoted as $\hat{Y}=[\hat{Y}^{1},\hat{Y}^{2},\ldots,\hat{Y}^{K}]\in\mathbb{R}^{K\times T_{f}\times 2}$. Our goal is to learn a generative model to parameterize the distribution $\mathcal{P}(Y|X,\mathcal{M})$.
III-B Prediction Model
We named our proposed model ASPILin, emphasizing agent selection, physical interactions, and linear time weighting. An overview of our proposed ASPILin is illustrated in Fig. 2.
III-B1 Interacting agent selection
We define the Euclidean distance between agent $i$ and the target agent at time $t$ as $d_{t}^{i}$. If $d_{t}^{i}$ falls below a manually defined selection threshold $\mathcal{D}$, it is considered that there is a potential for interaction between agent $i$ and the target agent.
In urban areas, vehicles operate within lanes while driving, which allows us to convert the problem of interaction between agents into a correlation problem between lanes. At time $t$, the lane to which agent $i$ belongs is defined as $l_{t}^{i}$. As time passes, the agent will move to another lane, which we refer to as future lane $l_{t+}^{i}$ ($t+\leq T_{f}$). Because the size of each scene is finite, if $l_{t}^{i}$ represents the last lane traversed by agent $n$ in the scene, then it is stipulated that $l_{t+}^{i}=l_{t}^{i}$. The future lane $l_{t+}^{i}$ can be easily captured for training data. However, a lane predictor must be built at the inference stage to predict $l_{t+}^{i}$. Intuitively, predicting the future lane can easily be converted into a classification problem. Nevertheless, this method is unsuitable for real-world applications as it only accommodates predictions in scenarios with training. For this purpose, we design an ultra-lightweight model Lin to forecast unimodal medium-to-high-precision trajectories for future moments, then map each timestep’s trajectory onto the respective lane. Lin is a simplified version of ASPILin, which excludes the interaction module. The entire process can be expressed as $\{l_{k}^{i}\}_{k=-T_{h}+1}^{0}=g(f_{\text{reg}}([\{x_{k}^{i}\}_{k=-T_{h}+1}^{0},\mathcal{M}]))$, where $g$ is the mapping from trajectory to lane. One challenge of using this regression method is that agents in a real-world scenario sometimes have different observation steps. Yet, the model must predict the trajectories of all agents to ensure the accuracy of the interacting agent selection. Therefore, during inference, we use a Constant Acceleration model $\text{CA}(\cdot)$ to pad the missing observation steps for agents. For an agent $i$ whose states $\{x_{k}^{i}\}_{k=t_{1}}^{t_{2}}$ ($-T_{h}+1<t_{1}\leq t_{2}<0$) are missing observation steps, the most recent state is used as the input for $\text{CA}(\cdot)$:
$\{x_{k}^{i}\}_{k=-T_{h}+1}^{t_{1}-1}=\text{CA}(x_{t_{1}}^{i}),$ (1)
$\{x_{k}^{i}\}_{k=t_{2}+1}^{0}=\text{CA}(x_{t_{2}}^{i}).$ (2)
Algorithm 1 Interacting agent selection algorithm
0:  Agents’ state $X_{t}\in\mathbb{R}^{N\times 7}$, Agents’ current and future lanes $L_{t}\in\mathbb{R}^{N\times 2}$, Selection threshold $\mathcal{D}$
0:  Index list $\mathcal{N}$
1:  Extract position $P_{t}$ and velocity $V_{t}$ from $X_{t}$;
2:  Initialize min-distance $d_{\text{SL}},d_{\text{FL}},d_{\text{FF}},d_{\text{ML}}\leftarrow\mathcal{D}$;
3:  Initialize neighbor index list $\mathcal{N}\leftarrow[0,0,0,0]$;
4:  for each agent $i\leftarrow 2$ to $N$ do
5:     Distance $d_{t}^{i}\leftarrow\|p_{t}^{i}-p_{t}^{1}\|_{2}$;
6:     Orientation $o_{t}^{i}\leftarrow(p_{t}^{i}-p_{t}^{1})\cdot v_{t}^{1}$;
7:     if $d_{t}^{i}<d_{\text{SL}}\ and\ l_{t}^{i}=l_{t}^{1}\ and\ o_{t}^{i}\geq 0$ then
8:        $d_{\text{SL}}\leftarrow d_{t}^{i}$;
9:        $\mathcal{N}[0]\leftarrow i$;
10:     else if $d_{t}^{i}<d_{\text{FL}}\ and\ l_{t}^{i}=l_{t+}^{1}\ and\ o_{t}^{i}\geq 0\ and\ l_{t+}^{1}\neq l_{t}^{1}$ then
11:        $d_{\text{FL}}\leftarrow d_{t}^{i}$;
12:        $\mathcal{N}[1]\leftarrow i$;
13:     else if $d_{t}^{i}<d_{\text{FF}}\ and\ l_{t}^{i}=l_{t+}^{1}\ and\ o_{t}^{i}<0\ and\ l_{t+}^{1}\neq l_{t}^{1}$ then
14:        $d_{\text{FF}}\leftarrow d_{t}^{i}$;
15:        $\mathcal{N}[2]\leftarrow i$;
16:     else if $d_{t}^{i}<d_{\text{ML}}\ and\ l_{t}^{1}=l_{t+}^{i}\ and\ o_{t}^{i}\geq 0\ and\ l_{t+}^{i}\neq l_{t}^{i}$ then
17:        $d_{\text{ML}}\leftarrow d_{t}^{i}$;
18:        $\mathcal{N}[3]\leftarrow i$;
19:     end if
20:  end for
21:  return  $\mathcal{N}$;
Then we select four different types of interacting agents as described in Alg. 1: Same lane and Leading (SL), Future lane and Leading (FL), Future lane and Following (FF), and Merging and Leading (ML). If the angle between an agent’s position relative to the target agent and the target agent’s heading does not exceed $90^{\circ}$, this agent is considered leading the target agent. Each type’s agent closest to the target agent is finally selected as its interacting agent.
III-B2 Interaction Representation
[…]
Similar to recent studies[2, 11, 7], we convert the coordinate system for the final interaction representation. Specifically, all states in $I$ are transformed into a relative coordinate system as $\tilde{I}$ with the target agent’s final observation point $p_{0}^{1}$ as the origin and the positive direction $h_{0}^{1}$ of the x-axis.
III-B3 Interaction Encoding
For the traditional multi-head dot-product attention mechanism, the interaction embedding $e_{int}$ between the target agent and all of its interacting agents can be denoted as:
$Z=\text{FC}(\tilde{I})=[Z^{1},(Z^{s})_{s\in\mathcal{S}}],$ (3)
$Q,K,V=W^{q}Z^{1},W^{k}(Z^{s})_{s\in\mathcal{S}},W^{v}(Z^{s})_{s\in\mathcal{S}},$ (4)
$Z^{\prime}=\text{MHA}(Q,K,V),$ (5)
$e_{int}=\text{LN}(\text{FFN}(\text{LN}(Z^{1}+Z^{\prime}))+Z),$ (6)
where $\text{FC}(\cdot)\in\mathbb{R}^{7T_{h}\times d_{z}}$ performs a direct linear transformation on the temporal dimension features, and $Q,K,V$ are obtained from three equivalent linear transformations ($W^{q},W^{k},W^{v}\in\mathbb{R}^{d_{z}\times d_{z^{\prime}}}$). Residual connection and layer normalization LN($\cdot$) are applied to the output $Z$ from MHA()($\cdot$)( ⋅ ) before input into the feed-forward module FFN($\cdot$), following the standard Transformer architecture.
We now introduce a physical method to calculate the correlation $a_{t}^{i_{1}}$ between agent $i_{1}$ and the target agent at time $t$. First, we use the $\text{CA}(\cdot)$ to estimate the time $\tau_{t}^{i_{1}}$ needed for the target agent and agent $i_{1}$ to reach the closest distance $d_{t+}^{i_{1}}$, expressed as:
[…]
$d_{t+}^{i_{1}}=\|\text{CA}(\tilde{x}_{t}^{1})-\text{CA}(\tilde{x}_{t}^{i_{1}})\|_{2}.$ (8)
Then, combining with the current distance $d_{t}^{i_{1}}$ to calculate the correlation coefficient :
$\overline{\tau}_{t}^{i_{1}}=\begin{cases}0&\text{if }\tau_{t}^{i_{1}}<0;\\ T&\text{if }\tau_{t}^{i_{1}}>T;\\ \tau_{t}^{i_{1}}&\text{otherwise},\end{cases}$ (9)
$c_{t}^{i_{1}}=\frac{d_{t}^{i_{1}}-d_{t+}^{i_{1}}+\epsilon}{d_{t}^{i_{1}}\text{exp}(\overline{\tau}_{t}^{i_{1}})}.$ (10)
Here, we set a lower bound for $\tau_{t}^{i_{1}}$ to ensure that the correlation coefficient is based solely on future action and an upper bound based on the assumption that agents will not stay in the scene for more than $T$ seconds, to prevent excessively high values in subsequent calculations. Eq. 10 indicates that the correlation between agents is negatively correlated with $d_{t}^{i_{1}}$ and $\tau_{t}^{i_{1}}$, and positively correlated with the difference in the distance they cover during $\tau_{t}^{i_{1}}$. Adding an extra term $\epsilon$ in the numerator prevents it from 0, thereby maintaining the effectiveness of the two factors in the denominator. Using $\text{exp}(\cdot)$ in the denominator serves to highlight the impact of $\tau_{t}^{i_{1}}$ and avoid the denominator being 0. If at time $t$ there is no SL type agent $i_{1}$, then $c_{t}^{i_{1}}$ is set to 0. Next, $c_{t}^{i_{1}}$ is normalized to a physical attention score:
$\alpha_{t}^{i_{1}}=\alpha_{t}^{\text{SL}}=\frac{2(T_{h}+t)c_{t}^{i_{1}}}{(1+T_{h})T_{h}\sum_{n=1}^{4}c_{t}^{i_{n}}},$ (11)
where $\alpha_{t}^{i_{1}}$ simultaneously incorporates the spatial attention of the target agent towards agent $i_{1}$ at time $t$ and the linear weighting of $t$ over the entire time scale $T_{h}$. Subsequently, we derive a weight matrix $\mathcal{A}\in\mathbb{R}^{4\times T_{h}}$:
$\mathcal{A}=\begin{bmatrix}\alpha_{-T_{h}+1}^{\text{SL}}&\cdots&\alpha_{0}^{\text{SL}}\\ \alpha_{-T_{h}+1}^{\text{FL}}&\cdots&\alpha_{0}^{\text{FL}}\\ \alpha_{-T_{h}+1}^{\text{FF}}&\cdots&\alpha_{0}^{\text{FF}}\\ \alpha_{-T_{h}+1}^{\text{ML}}&\cdots&\alpha_{0}^{\text{ML}}\end{bmatrix}=\begin{bmatrix}\mathcal{A}^{\text{SL}}\\ \mathcal{A}^{\text{FL}}\\ \mathcal{A}^{\text{FF}}\\ \mathcal{A}^{\text{ML}}\end{bmatrix}.$ (12)
Ultimately, $e_{int}$ is obtained by applying residual connection and layer norm:
$Z=\text{FC}_{z}(\tilde{I})=[Z^{1},(Z^{s})_{s\in\mathcal{S}}],$ (13)
$Z^{\prime}=\text{FC}_{z^{\prime}}\left(Z^{1}+\sum_{s\in\mathcal{S}}\mathcal{A}^{s}\circ Z^{s}\right),$ (14)
$e_{int}=\text{FFN-LN}(\text{FFN}(\text{LN}(Z^{\prime})))+Z^{\prime},$ (15)
where $\text{FC}_{z}(\cdot)\in\mathbb{R}^{7\times d_{z}}$ and $\text{FC}_{z^{\prime}}(\cdot)\in\mathbb{R}^{d_{z}T_{h}\times d_{z^{\prime}}}$ perform linear transformations on spatial and temporal features respectively, $\circ$ denotes the Hadamard product. Inspired by NormFormer[24], we additionally use a LN placed after the FNN()($\cdot$)( ⋅ ) but before the residual connection on top of the Pre-LN configuration, referred to as FFN-LN()($\cdot$)( ⋅ ), which helps enhance training stability.
III-B4 Spatiotemporal Encoder
1D Convolutional Neural Networks Conv1D($\cdot$) and Gated Recurrent Units GRU($\cdot$) are used to capture the spatial and temporal dependencies of the target agent’s historical states in the relative coordinate system, ultimately resulting in a spatiotemporal embedding $e_{st}$:
$e_{st}=\text{GRU}(\text{Conv1D}(\{\tilde{x}_{k}^{1}\}_{k=-T_{h}+1}^{0})).$ (16)
We use three equivalent but independent spatiotemporal encoders corresponding to decoders to obtain the mean $\mu$, variance $\sigma$, and samples $z$ used for reparameterization respectively. The corresponding embeddings are labeled as $e_{st}^{\mu}$, $e_{st}^{\sigma}$, and $e_{st}^{z}$.
III-B5 Map Encoder
The map selector from HEAT-I-R[2] is utilized as our map encoder. The entire process is as follows:
$\mathcal{M}^{\prime}=\text{MLP}(\text{CNNs}(\mathcal{M})),$ (17)
$x^{\prime}=\text{FC}_{x^{\prime}}(x_{0}^{1}),$ (18)
$e_{map}=\text{Sigmoid}(\text{FC}_{m}([\mathcal{M}^{\prime},x^{\prime}]))\circ\mathcal{M}^{\prime}.$ (19)
CNNs($\cdot$) and MLP($\cdot$) are utilized to extract features from the map and perform a linear transformation $\text{FC}_{x^{\prime}}(\cdot)\in\mathbb{R}^{7\times d_{x^{\prime}}}$ of the target agent’s state at the current time from global coordinates into a higher-dimensional space. Then, we fuse the features of $x^{\prime}$ and $\mathcal{M}^{\prime}$ ($\text{FC}_{m}(\cdot)\in\mathbb{R}^{(d_{x^{\prime}}+d_{\mathcal{M}^{\prime}})\times d_{\mathcal{M}^{\prime}}}$), pass them through Sigmoid($\cdot$) to scale all feature values between 0 and 1, and finally perform element-wise production with $\mathcal{M}^{\prime}$ to obtain the map embedding $e_{map}$.
III-B6 Decoder and Reparameterization
The mean and variance of the future trajectory of the target agent are forecasted at first, denoted as:
$\mu=\text{MLP}_{\mu}([e_{st}^{\mu},e_{int},e_{map}]),$ (20)
$\sigma=\text{MLP}_{\sigma}([e_{st}^{\sigma},e_{int},e_{map}]),$ (21)
where $\mu\in\mathbb{R}^{T_{f}\times 2}$ represents a set containing the means of the trajectory for each future timestep, and $\sigma\in\mathbb{R}$ is a single value. Next, we perform a set of linear transformations on the predicted $\sigma$ and use it for sample prediction:
$\sigma^{\prime}=\text{MLP}_{\sigma^{\prime}}(\sigma),$ (22)
$z=\text{MLP}_{z}([e_{st}^{z},e_{int},e_{map},\sigma^{\prime}]),$ (23)
where $z\in\mathbb{R}^{K\times T_{f}\times 2}$ represents the $K$ samples for each future timestep. The predicted trajectory $\hat{Y}$ is ultimately derived from a reparameterization formula:
$\hat{Y}=\mu+\sigma\times z.$ (24)
III-C Training Objective
The loss function inspired by Leapfog[6] used to optimize both ASPILin and Lin is expressed as:
$\mathcal{L}=\mathcal{L}_{\text{distance}}+\lambda\mathcal{L}_{\text{diversity}},$ (25)
$\mathcal{L}_{\text{distance}}=\frac{1}{T_{f}}{\min}_{k=1}^{K}\sum_{t=1}^{T_{f}}\|\hat{y}_{t}^{k}-y_{t}\|_{2},$ (26)
$\mathcal{L}_{\text{diversity}}=\frac{\sum_{k=1}^{K}\sum_{t=1}^{T_{f}}\|\hat{y}_{t}^{k}-y_{t}\|_{2}}{\sigma^{2}KT_{f}}+\log\sigma^{2}.$ (27)
We apply a winner-takes-all strategy to $\mathcal{L}_{\text{distance}}$, performing backpropagation only on the mode with the smallest deviation from the ground truth to prevent mode collapse. $\mathcal{L}_{\text{diversity}}$ with $\lambda=0.02$ is designed to enhance the diversity of forecasted trajectories. The first component aligns variance with scene complexity, while the second component acts as a regularization factor, ensuring high variance across all predictions.
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
We train and evaluate our model on three interaction-rich datasets: the INTERACTION dataset[14], the highD dataset[15], and the CitySim dataset[16]. INTERACTION contains 398,409 and 107,269 sequences for ASPILin’s training and validation and 413,548 and 111,493 sequences for Lin’s training and validation. Each sequence is sampled at 10Hz, and the task is to use the past 1 second of sequence data to predict the next 3 seconds. In the case of highD, we split the dataset into 2,530,166 training sequences, 180727 validation sequences, and 361,454 test sequences (7:1:2 ratio) and downsample the trajectories to 5Hz, following the data processing steps in PiP[25]. The prediction task involves using the past 3 seconds to predict the next 5 seconds. For CitySim, we use data from two no-signal scenarios, Intersection B and Roundabout A, with the training and validation split in an 8:2 ratio. The trajectory sampling rate for both scenarios is 30Hz, with the task being to predict the future 6 seconds trajectory based on the past 2 seconds. ASPILin and Lin use the same data for training and validation, including 61,185 training sequences and 15,164 validation sequences.
IV-A2 Metrics
For INTERACTION and CitySim, we forecast future trajectories for $K=6$ modes and evaluate the model’s performance using Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE), denoted as:
$\text{minADE}_{K}=\frac{1}{T_{f}}{\min}_{k=1}^{K}\sum_{t=1}^{T_{f}}\|\hat{y}_{t}^{k}-y_{t}\|_{2},$ (28)
$\text{minFDE}_{K}={\min}_{k=1}^{K}\|\hat{y}_{T_{f}}^{k}-y_{T_{f}}\|_{2}.$ (29)
For highD, we predict a deterministic unimodal trajectory and evaluate the model using Root Mean Square Error (RMSE), expressed as:
IV-A3 Implementation Details
$\text{RMSE}=\sqrt{\frac{1}{NT_{f}}\sum_{i=1}^{N}\sum_{t=1}^{T_{f}}\|\hat{y}_{t,i}-y_{t,i}\|_{2}^{2}},$ (30)
where $N$ represents the total number of samples.
The interaction module’s range threshold $\mathcal{D}$ is set to 30/200/45 meters for the INTERACTION/highD/CitySim dataset. Upper bound $T$ is set to 30, and extra term $\epsilon$ is set to 1. In our proposed physics-related method, $d_{z}$ and $d_{z^{\prime}}$ are configured as 32 and 256, respectively, and the feed-forward module has a dimension of 256. For the spatiotemporal encoder, the Conv1D kernel size is set to 3, the output channels to 32, and the GRU’s hidden layer dimension to 256. For the map encoder, we use the same settings as HEAT-I-R[2], where $\mathcal{M}$ is a $400\times 250$ grayscale map for each scene. For the three decoders, the hidden layers of $\text{MLP}_{\mu}(\cdot)$ and $\text{MLP}_{\sigma}(\cdot)$ are set to (256, 128), and the hidden layer of $\text{MLP}_{z}(\cdot)$ is set to (1024, 1024). The hidden layer and output dimension of linear transformations $\text{MLP}_{\sigma^{\prime}}(\cdot)$ are set to (8, 16) and 32, respectively.
ASPILin (3.5M/2.4M/5.8M parameters for the INTERACTION/highD/CitySim dataset) and Lin (2.5M/0.8M/2.9M parameters for the INTERACTION/highD/CitySim dataset) are trained on a single RTX-4090. We use AdamW as the optimizer, with a cosine annealing scheduler[26]. The initial settings for the learning rate, batch size, and training epochs are 1e-3, 64/128/32 for the INTERACTION/highD/CitySim dataset, and 40, respectively.
TABLE I: Comparison with models on the INTERACTION validation set
Model Conference minADE6$\downarrow$ minFDE6$\downarrow$
HEAT-I-R[2] * TITS 2022 0.19 0.66
ITRA[3] ITSC 2021 0.17 0.49
GOHOME[13] ICRA 2022 - 0.45
joint-StarNet[1] IV 2022 0.13 0.38
DiPA[11] RAL 2023 0.11 0.34
MB-SS-ASP[27] IV 2023 0.10 0.30
SAN[28] IV 2022 0.10 0.29
GMM-CVAE[23] arXiv 2023 0.09 0.28
Lin * - 0.18 0.67
ASPILin - 0.08 0.27
  • *Model that only performs unimodal prediction.
TABLE II: Comparison with models on the highD test set
[…]
1s 2s 3s 4s 5s
CV - 0.11 0.35 0.73 1.24 1.86
PiP[25] ECCV 2020 0.17 0.52 1.05 1.76 2.63
MHA-LSTM[29] TIV 2021 0.06 0.09 0.24 0.59 1.18
MMnTP[30] RAL 2023 0.19 0.38 0.62 0.95 1.39
POVL[31] ITSC 2023 0.12 0.18 0.22 0.53 1.15
VVF-TP[32] ITSC 2023 0.12 0.24 0.41 0.66 0.98
iNATran[33] TITS 2022 0.04 0.05 0.21 0.54 1.10
BAT[34] AAAI 2023 0.08 0.14 0.20 0.44 0.62
Lin - 0.05 0.06 0.11 0.27 0.54
ASPILin - 0.05 0.05 0.08 0.21 0.42
TABLE III: Ablation experiments for each component of the interaction module
Variant Agent Selection Time Windows Interactions Encode INTERACTION CitySim
four lane-related four closest all current physics-related attention minADE6$\downarrow$ minFDE6$\downarrow$ minADE6$\downarrow$ minFDE6$\downarrow$
1 0.119 0.393 0.999 2.223
2 0.117 0.392 1.010 2.253
3 0.096 0.321 0.982 2.178
4 0.097 0.331 0.973 2.164
5 0.091 0.314 0.965 2.137
6 0.088 0.295 0.959 2.124
7 0.079 0.266 0.954 2.095
IV-B Comparison with State-of-the-art
The results on INTERACTION shown in Tab. I indicate that ASPILin achieves SOTA performance. Moreover, as a lightweight model that does not account for interactions, Lin still achieves the semblable performance as models from the past 2-3 years, demonstrating the feasibility of our agent selection approach. The comparison results on highD are shown in Tab. II. Similar to some work[30, 31], we implemented a Constant Velocity (CV) model as a reference baseline. Interestingly, the performance after 1s exhibited by Lin exceeds that of SOTA methods, attributable to its superior CVAE architecture and loss functions. By comparison, ASPILin significantly reduces prediction error after 2s, highlighting the exceptional performance of our interaction modeling for long-term prediction.
IV-C Ablation Studies
IV-C1 Components of the Interaction Module
The ablation experiments for each component of the interaction module are shown in Tab III. The baseline configuration selects all agents within $\mathcal{D}$ at time $t=0$ as interacting agents and encodes their interactions using Transformer. Additionally, we introduced an extra simple agent selection method, which selects the four closest agents to verify the effectiveness of our method. An intuitive conclusion is that merely setting an upper limit on the number of interaction agents does not enhance model performance and may even reduce it. This is reasonable, as inappropriately narrowing the selection range will likely exclude genuinely interacting agents. However, switching the time window from current to all results in enhanced model performance, demonstrated across all three comparison sets (variants 2 and 3, 4 and 5, 6 and 7). Another interesting observation is that variants 2 and 3 are more significant than the other two groups. That’s because the four closest agents in interaction-dense scenarios may change at any moment, whereas four intent-matching agents are less likely to change over time. From comparisons in two other groups (variants 2 and 4, 3 and 5), we conclude that refining agent selection through lane usage can improve model performance, which provides valuable insights for future research. The last two comparisons (variants 4 and 6, 5 and 7) demonstrate that integrating physical interaction encoding is viable and advantageous, increasing the model’s interpretability.
IV-C2 Four Types of Interacting Agents
TABLE IV: Ablation experiments for four types of interacting agents
SL FL FF ML INTERACTION CitySim
minADE6$\downarrow$ minFDE6$\downarrow$ minADE6$\downarrow$ minFDE6$\downarrow$
0.083 0.282 1.013 2.235
0.080 0.269 1.025 2.307
0.090 0.299 1.007 2.243
0.085 0.271 0.962 2.104
0.079 0.266 0.954 2.095
According to the results in Tab. IV, excluding any category of interacting agents results in some level of decline in model performance. Models excluding FF and FL, achieve the poorest performance on INTERACTION and CitySim, indicating that agents on the target agent’s future lane have a more significant impact on the target agent than other agents. Moreover, on the CitySim validation set, the prediction task is more sensitive to changes in interaction agents because of its 6-second prediction horizon.
Figure 3: Qualitative results on INTERACTION. Purple rectangles represent vehicles, while green circles denote pedestrians or cyclists. Past trajectories are shown with black lines, predicted future trajectories with red lines, and ground truth trajectories with purple lines, with endpoints marked distinctively.
IV-C3 Different Lane Predictors
Despite our clarification that the lane predictor based on classification is not practical for real-world applications, we implemented this approach to examine how the accuracy of future lane predictions affects model performance by a simple Random Forest (RF). The detailed experimental results are shown in Tab. V. Using raw data undoubtedly achieves the best performance. Interestingly, the classification- and regression-based models showed starkly different performances on the two datasets because the prediction tasks on the CitySim dataset are considerably more complex than those on INTERACTION, especially since interactions are not considered in both. What is not displayed in the table is that Lin’s ADE and FDE on the CitySim dataset are 2.809 and 7.785, respectively. Nevertheless, it maintains a high lane prediction accuracy, demonstrating that our comprehensive agent selection strategy is effective even for long-term prediction tasks with a simple model. Additionally, it can be observed that different methods have reduced the discrepancies in lane prediction in their final forecasts. This is because our agent selection is not entirely based on future lanes, and predicting an incorrect future lane may still lead to the same outcome due to other mismatched conditions.
TABLE V: Ablation experiments for different lane predictors
Dataset Model ACC(%$\downarrow$ minADE6$\downarrow$ minFDE6$\downarrow$
INTERACTION Clf. (RF) 88.5 0.080 0.270
Reg. (Lin) 97.9 0.079 0.266
Raw Data 100 0.078 0.261
CitySim Clf. (RF) 93.6 0.951 2.089
Reg. (Lin) 89.5 0.954 2.095
Raw Data 100 0.946 2.073
IV-D Qualitative Results
We present the qualitative results of ASPILin on the INTERACTION dataset. As illustrated in Fig. 3, our model can predict accurate, multi-modal vehicle trajectories in complex scenarios. Interestingly, ASPILin struggles more with predicting the precise trajectories of vehicles that are stationary or nearly so, which points to potential areas for future enhancements.
IV-E Inference Latency
TABLE VI: Inference latency for three datasets
Dataset Model LP AS TP Total(ms)$\downarrow$
INTERACTION baseline - 22.29 0.51 23.20
ASPILin 1.00 8.52 0.24 9.76
CitySim baseline - 6.66 0.79 7.45
ASPILin 0.99 6.53 0.26 7.78
We evaluate the inference latency of the entire prediction process, including Lane Prediction (LP), Agent Selection (AS), and Trajectory Prediction (TP) on INTERACTION and CitySim and compare it with the baseline (i.e., variant 1 in Tab. III). AS consumes most computational resources due to extensive data processing and conditional filtering. Even though ASPILin employs more criteria for meticulous agent selection, it remains more efficient than the baseline. In the INTERACTION dataset, up to 25 agents can be within 30 meters of the target agent, while in CitySim, there are a maximum of only 5 agents within 45 meters. This explains the discrepancies between ASPILin and the baseline in the two datasets. Results show that ASPILin possesses high inference efficiency, particularly in interaction-rich scenes. While its efficiency slightly underperforms the baseline in scenarios with low agent density, this is entirely acceptable.
V Conclusion
In this work, we explored the possibility of enhancing interaction for trajectory prediction from two perspectives: (i) a lane-related method for a more detailed selection of interacting agents, and (ii) a physically-related interaction encoding method. We designed a model named ASPILin and conducted experiments on popular datasets. The results indicate that our approach significantly enhances interaction modeling and positively affects trajectory prediction, offering substantially increased interpretability over earlier methods.